# Exploring Holomorphic Retracts

### v6

@@ Conclusion
The foregoing arguments – again using theorem 1.3 at the outset to reduce considerations from the generality of holomorphic retracts to linear ones – can be adapted to determine precisely the retracts through the origin of other ‘simpler’ examples of non-convex (but pseudoconvex) bounded Reinhardt domains such as $\mathbb{B}\cap H^{2}_{1/4}$ or $B_{p}\cap H^{2}_{1/4}$, where $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/4\}$ and where $B_{p}$ denotes the standard $\ell^{p}$-ball with $1\leq p<\infty$. The case $p=\infty$ is different as in that case (i.e., for $\Delta^{2}\cap H^{2}_{1/4}$) there are non-linear retracts through the origin and therefore, discussed separately below. The reason we chose to deal with $D_{h}$ extensively is not because it fails to be Reinhardt (though that is an added virtue) but rather because it fails to be non-convex in a maximal manner: note that there does not exist any bounded domain which is non-convex at all boundary points, as follows by standard considerations involving farthest point(s) from a fixed point in the interior and appealing to the Hahn – Banach theorem; for a balanced domain the next best thing to ask for is one that is non-convex at an open dense subset of its boundary or, that the set of convexity points is of measure zero. Both of these conditions are fulfilled by $D_{h}$. Despite its simplicity, one can still ask natural questions about the choice of $D_{h}$; without elaborating on them all, let us just mention one: why not consider $D_{q}\cap H^{2}_{1/4}$ (where $D_{q}$ denotes the ‘$\ell_{q}$-ball’ for $10<q<$ 1 as before), as this is also a maximally non-convex domain? Our answer to this is that though this is easier to analyse owing to it being Reinhardt, it turns out that through every non-smooth point of $D_{q}\cap H^{2}_{1/4}$ there is a linear retract and is a convex boundary point whereas this is not the case for $D_{h}$: not every non-smooth point of $\partial D_{h}$ is a point of convexity, thereby is an even better example of a maximally non-convex domain. Moreover, through almost every point $\partial D_{h}$, there passes analytic discs which lie within $\partial D_{h}$ making it at the other extreme compared to convex domains of finite type which are by far, well-studied. Let us conclude this remark by finishing off what we began it with namely, retracts of $B_{p}\cap H^{2}_{1/4}$ for $1\leq p<\infty$ which provide simple examples of balanced domains which are neither convex nor maximally non-convex i.e., lie somewhere in-between these extremes (the role of intersection with $H^{2}_{1/4}$ is mainly to create this intermediate possibility by destroying the convexity of $B_{p}$ and could be replaced with other domains). The retracts through the origin of such domains are precisely those points of $\partial B_{p}$ which remain after the intersection with $H^{2}_{1/4}$ is performed. Let us also draw attention to the fact that the case $p=1$ is allowed whereas $p=\infty$ is not because there are plenty of non-linear retracts (through the origin) in the polydisc.

### v7

@@ Conclusion
The foregoing arguments – again using proposition 1.3 at the outset to reduce considerations from the generality of holomorphic retracts to linear ones – can be adapted to determine precisely the retracts through the origin of other ‘simpler’ examples of non-convex (but pseudoconvex) bounded Reinhardt domains such as $\mathbb{B}\cap H^{2}_{1/4}$ or $B_{p}\cap H^{2}_{1/4}$, where $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}\;:\;|zw|<1/4\}$ and where $B_{p}$ denotes the standard $\ell^{p}$-ball with $1\leq p<\infty$. The case $p=\infty$ is different as in that case (i.e., for $\Delta^{2}\cap H^{2}_{1/4}$) there are non-linear retracts through the origin and therefore, discussed separately below. The reason we chose to deal with $D_{h}$ extensively is not because it fails to be Reinhardt (though that is an added virtue) but rather because it fails to be non-convex in a maximal manner: note that there does not exist any bounded domain which is non-convex at all boundary points, as follows by standard considerations involving farthest point(s) from a fixed point in the interior and appealing to the Hahn – Banach theorem; for a balanced domain the next best thing to ask for is one that is non-convex at an open dense subset of its boundary or, that the set of convexity points is of measure zero. Both of these conditions are fulfilled by $D_{h}$. Despite its simplicity, one can still ask natural questions about the choice of $D_{h}$; without elaborating on them all, let us just mention one: why not consider $D_{q}\cap H^{2}_{1/4}$ (where $D_{q}$ denotes the $\ell_{q}$-ball’ for $0<q<1$ as before), as this is also a maximally non-convex domain? Our answer to this is that though this is easier to analyse owing to it being Reinhardt, it turns out that through every non-smooth point of $D_{q}\cap H^{2}_{1/4}$ there is a linear retract and is a convex boundary point whereas this is not the case for $D_{h}$: not every non-smooth point of $\partial D_{h}$ is a point of convexity, thereby is an even better example of a maximally non-convex domain. Moreover, through almost every point $\partial D_{h}$, there passes analytic discs which lie within $\partial D_{h}$ making it at the other extreme compared to convex domains of finite type which are by far, well-studied. Let us conclude this remark by finishing off what we began it with namely, retracts of $B_{p}\cap H^{2}_{1/4}$ for $1\leq p<\infty$ which provide simple examples of balanced domains which are neither convex nor maximally non-convex i.e., lie somewhere in-between these extremes (the role of intersection with $H^{2}_{1/4}$ is mainly to create this intermediate possibility by destroying the convexity of $B_{p}$ and could be replaced with other domains). The retracts through the origin of such domains are precisely those points of $\partial B_{p}$ which remain after the intersection with $H^{2}_{1/4}$ is performed. Let us also draw attention to the fact that the case $p=1$ is allowed whereas $p=\infty$ is not because there are plenty of non-linear retracts (through the origin) in the polydisc.